OpenStudy (anonymous):

how do i find a zero vector when given X:=(1,x)

OpenStudy (zehanz):

Do you mean a vector perpendicular to this one, so the inproduct of the two will be zero? A zero vector would just be (0,0).

OpenStudy (anonymous):

if you mean perpendicular/orthogonal, then it would be \((-1,~x-x)\)

OpenStudy (zehanz):

@yummydum: I think you've made a mistake. If theinproduct is to be zero, switch the numbers and negate of them: (1,x) • (x, -1) = 1*x + x*(-1) = x - x =0. So (x, -1) would be a good choice, or (-x,1).

OpenStudy (anonymous):

oops! sorry yes, i was thinking of something else

OpenStudy (zehanz):

Hey...@yummydum! Stay focused! ;)

OpenStudy (anonymous):

lol, I'm tryin